# Multi-Preference Lambda-weighted Listwise DPO for Small-Scale Model Alignment

### v5

@@ A.1 Deriving the Final Training Objective
$\mathcal{L}^{\text{DPO}}(\theta)=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\exp(r_{\theta}(x,y_{i}))}{\sum_{j=1}^{N}\exp(r_{\theta}(x,y_{j}))}\right)\right]$ (20)
@@
$\mathcal{L}^{\text{DPO}}(\theta)&=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}Z(x)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}Z(x)^{\beta}}\right)\right]\\ &=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}\right)\right]\end{aligned}$
@@
$\mathcal{L}^{(k)}(\theta)=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p_{i}^{*(k)}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}\right)\right]$ (21)
@@
$\mathcal{L}_{\text{final}}(\theta)=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\mathcal{L}^{\lambda}(\theta)\right]=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\sum_{k=1}^{m}\lambda_{k}\cdot\mathcal{L}^{(k)}(\theta)\right]$ (23)
@@ A.2 From Final Objective to Lambda-DPO Form
$\mathcal{L}_{\text{final}}(\theta)$ $=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\sum_{k=1}^{m}\lambda_{k}\cdot\mathbb{E}_{(x,\{y_{i}\})}\left[-\sum_{i=1}^{N}p_{i}^{*(k)}\log P_{\theta}(y_{i}|x)\right]\right]$ (25)
@@
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (28)
@@
A.3 Gradient of Lambda-DPO Loss
@@
$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}$ $=-\mathbb{E}_{x,\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\nabla_{\theta}\log P_{\theta}(y_{i}\mid x)\right]$ (29)
$=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$
@@
Appendix B B Experiment Setup
B.1 Experimental Objectives and Comparison Targets
The primary goal of our experiments is to evaluate the performance of the proposed $\lambda$-DPO algorithm compared to the standard DPO method. Specifically, we assess their effectiveness across different model architectures and training settings using public preference datasets and standard evaluation benchmarks.
We consider two levels of comparison:
Model comparison.
We experiment with two strong open-source language models—LLaMA and Qwen. These models differ in architecture, pretraining corpora, and tokenization schemes, providing a robust testbed for evaluating generalizability.
Method comparison.
We compare DPO and $\lambda$-DPO. For each method-model pair, we perform independent hyperparameter tuning (e.g., learning rate, batch size, $\beta$) to ensure comparisons are fair and reflect the best achievable performance under comparable resource constraints.
B.2 Training Settings
Inspired by the SimPO baseline, we adopt a unified training pipeline:
  • Finetuned-Base: The model is first SFT-trained on UltraChat-200k, then optimized on UltraFeedback via DPO or $\lambda$-DPO to reflect realistic alignment scenarios.
B.3 Evaluation Benchmarks and Metrics
We evaluate model performance using the lm-eval-harness (Gao et al. 2024) framework, which reports accuracy on diverse benchmarks covering reasoning, factual correctness, and commonsense understanding.
Specifically, we report average accuracy across the following six established datasets:
  • MMLU: 57 subjects, 15K questions assessing academic knowledge (Hendrycks et al. 2020).
  • ARC: 1.1K grade-school science questions requiring reasoning (Clark et al. 2018).
  • HellaSwag: 10K commonsense sentence completions with distractors (Zellers et al. 2019).
  • TruthfulQA: 817 questions testing factuality and hallucination (Lin, Hilton, and Evans 2021).
  • Winograd: 273 examples probing coreference and syntax (Levesque, Davis, and Morgenstern 2012).
  • GSM8K: 8.5K arithmetic reasoning problems (Cobbe et al. 2021).
B.4 Ablation: Impact of Lambda Weighting
To better understand the impact of the $\lambda$ vector in $\lambda$-DPO, we conduct ablation experiments focusing on how different preference weightings affect alignment performance.
Specifically, we compare two $\lambda$ schedule variants within the listwise $\lambda$-DPO framework:
  • Uniform $\lambda$: Equal weights across all dimensions ($\lambda=(1/m,\dots,1/m)$), simulating balanced multi-objective aggregation.
  • Unimodal $\lambda$: One-hot vectors (e.g., $(1,0,\dots)$) isolating single-dimension optimization without interpolation.
This comparison helps isolate the contribution of $\lambda$-based preference interpolation within the listwise setting.
Appendix C C Accuracy Comparisons on UltraFeedback Benchmarks with 1B model
We provide benchmark accuracy comparisons between DPO and $\lambda$-DPO across multiple language models on the UltraFeedback dataset. Performance is reported across six tasks: mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k.
C.1 Qwen-1.8 (Base)
Figure 7 shows that $\lambda$-DPO achieves modest improvements over standard DPO on most tasks using the Qwen-1.8 base model.
Figure 7: Accuracy comparison between DPO and $\lambda$-DPO on Qwen-1.8 (base).
C.2 Qwen-1.8 Chat
Figure 8 presents results on Qwen-1.8 Chat, where $\lambda$-DPO consistently improves accuracy, especially on reasoning-heavy benchmarks like arc and hellaswag.
Figure 8: Accuracy comparison between DPO and $\lambda$-DPO on Qwen-1.8 Chat.
C.3 LLaMA 3.2 1B (Base)
As shown in Figure 9, $\lambda$-DPO outperforms DPO slightly on LLaMA 3.2 1B base across four out of six tasks, demonstrating generalization beyond the Qwen series.
Figure 9: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B base model.
C.4 LLaMA 3.2 1B (Instruct)
Figure 10 illustrates results for the Instruct-tuned variant of LLaMA 3.2 1B. The performance gap between $\lambda$-DPO and DPO is generally preserved, indicating robustness to instruction tuning.
Figure 10: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B Instruct.
C.5 Unimodal Ablation
We visualize the performance of four unimodal $\lambda$-DPO models, each trained using a one-hot preference vector that activates only one of the four human preference dimensions: helpfulness, honesty, coherence, and fluency. Accuracy is reported across six standard benchmarks: mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k.
Figure 11: Accuracy comparison of four unimodal $\lambda$-DPO models, each trained with a single active preference dimension. All models perform similarly across benchmarks, with slight advantages for helpfulness and honesty on average. This suggests that no single preference dominates, and multi-dimensional interpolation may improve overall alignment.
Appendix D D Performance-Driven Lambda Scheduling via Polynomial Modeling
@@
D.1 Polynomial Feature Mapping
@@ D.2 Polynomial Regression
D.3 Grid Construction over the Simplex
@@
D.4 Softmax-Based Scheduling Distribution
@@
D.5 Empirical Instantiation with Five Training Points
@@
$\lambda^{(1)}$ $=[1,0,0,0],\quad y^{(1)}=0.4593$
$\lambda^{(2)}$ $=[0,1,0,0],\quad y^{(2)}=0.4598$
$\lambda^{(3)}$ $=[0,0,1,0],\quad y^{(3)}=0.4543$
$\lambda^{(4)}$ $=[0,0,0,1],\quad y^{(4)}=0.4548$
$\lambda^{(5)}$ $=[0.25,0.25,0.25,0.25],\quad y^{(5)}=0.4612$
@@
$f(\lambda)$ $=0.35+0.09\lambda_{1}+0.09\lambda_{2}+0.09\lambda_{3}+0.09\lambda_{4}$
$+0.02\lambda_{1}^{2}+0.04\lambda_{1}\lambda_{2}+0.03\lambda_{1}\lambda_{3}+0.02\lambda_{1}\lambda_{4}$
$+0.02\lambda_{2}^{2}+0.02\lambda_{2}\lambda_{3}+0.02\lambda_{2}\lambda_{4}$
$+0.01\lambda_{3}^{2}+0.00\lambda_{3}\lambda_{4}+0.01\lambda_{4}^{2}$
D.6 Lambda Sampling from Learned Scheduler
We sample $k=10$ $\lambda$ vectors from Dirichlet(1,1,1,1)$1$(1,1,1,1)( 1 , 1 , 1 , 1 ) and compute predicted scores $f(\lambda^{(j)})$, converting them via softmax ($\tau=100$):
@@ D.6 Lambda Sampling from Learned Scheduler
D.7 Scheduler vs. Uniform Comparison
We compare our learned scheduler with a uniform $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ baseline using LLaMA 3.2B Instruct. The learned scheduler achieves slightly better average accuracy (0.462 vs. 0.461), and offers regularization via stochastic $\lambda$ sampling, enhancing robustness.
Figure 12: Comparison of learned scheduler vs. fixed uniform $\lambda$ across six alignment benchmarks. The learned scheduler yields improved accuracy and diversity.

### v6

@@ A.1 Deriving the Final Training Objective
$\mathcal{L}^{\text{DPO}}(\theta)=$ $\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\bigg[-\sum_{i=1}^{N}p^{*}_{i}$
$\cdot\log\left(\frac{\exp(r_{\theta}(x,y_{i}))}{\sum_{j=1}^{N}\exp(r_{\theta}(x,y_{j}))}\right)\bigg]$ (17)
@@
$\begin{aligned} &\mathcal{L}^{\text{DPO}}(\theta)\\ =&\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}Z(x)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}Z(x)^{\beta}}\right)\right]\\ =&\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}\right)\right]\end{aligned}$
@@
$\mathcal{L}^{(k)}(\theta)=$ $\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\bigg[-\sum_{i=1}^{N}p_{i}^{*(k)}$
$\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}\right)\bigg]$ (18)
@@
$\mathcal{L}_{\text{final}}(\theta)$ $=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\mathcal{L}^{\lambda}(\theta)\right]$ (20)
$=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\sum_{k=1}^{m}\lambda_{k}\cdot\mathcal{L}^{(k)}(\theta)\right]$ (21)
@@ A.2 From Final Objective to Lambda-DPO Form
$\mathcal{L}_{\text{final}}(\theta)$ $=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\Bigg\{\sum_{k=1}^{m}\lambda_{k}$
$\cdot\mathbb{E}_{(x,\{y_{i}\})}\left[-\sum_{i=1}^{N}p_{i}^{*(k)}\log P_{\theta}(y_{i}|x)\right]\Bigg\}$ (23)
@@
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)$
$=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (26)
@@
A.3 A.3 Gradient of Lambda-DPO Loss
@@
$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}$
$=-\mathbb{E}_{x,\lambda}$ $\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\nabla_{\theta}\log P_{\theta}(y_{i}\mid x)\right]$ (27)
$=-\mathbb{E}_{x,\lambda}$ $\bigg[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)$
$\qquad\quad\cdot\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\bigg]$ (28)
@@
Appendix B Benchmark Statistics
We report average accuracy across the following six established datasets:
  • MMLU: 57 subjects, 15K questions assessing academic knowledge Hendrycks et al. (2021).
  • ARC: 1.1K grade-school science questions requiring reasoning Clark and others (2018).
  • HellaSwag: 10K commonsense sentence completions with distractors Zellers et al. (2019).
  • TruthfulQA: 817 questions testing factuality and hallucination Lin et al. (2022).
  • Winograd: 273 examples probing coreference and syntax Levesque and others (2012).
  • MathQA: 37k Math reasoning problems Amini et al. (2019).
Appendix C Performance-Driven Lambda Scheduling via Polynomial Modeling
@@
C.1 Polynomial Feature Mapping
@@ C.2 Polynomial Regression
C.3 Grid Construction over the Simplex
@@
C.4 Softmax-Based Scheduling Distribution
@@
C.5 Empirical Instantiation with Five Training Points
@@
$\lambda^{(1)}$ $=[1,0,0,0],\quad y^{(1)}=0.4563$
$\lambda^{(2)}$ $=[0,1,0,0],\quad y^{(2)}=0.4561$
$\lambda^{(3)}$ $=[0,0,1,0],\quad y^{(3)}=0.4578$
$\lambda^{(4)}$ $=[0,0,0,1],\quad y^{(4)}=0.4553$
$\lambda^{(5)}$ $=[0.25,0.25,0.25,0.25],\quad y^{(5)}=0.4623$
@@
$f(\lambda)$ $=0.36+0.09\lambda_{1}+0.09\lambda_{2}+0.09\lambda_{3}+0.09\lambda_{4}$
$+0.0061\lambda_{1}^{2}+0.0280\lambda_{1}\lambda_{2}+0.0280\lambda_{1}\lambda_{3}$
$+0.0280\lambda_{1}\lambda_{4}+0.0060\lambda_{2}^{2}+0.0280\lambda_{2}\lambda_{3}$
$+0.0280\lambda_{2}\lambda_{4}+0.0068\lambda_{3}^{2}$
$+0.0280\lambda_{3}\lambda_{4}+0.0056\lambda_{4}^{2}.$
C.6 Lambda Sampling from Learned Scheduler
We sample $k=10$ $\lambda$ vectors from Dirichlet(1,1,1,1)(1,1,1,1) and compute predicted scores $f(\lambda^{(j)})$, converting them via softmax ($\tau=100$):
@@ C.6 Lambda Sampling from Learned Scheduler
C.7 Scheduler vs. Uniform Comparison
We compare our learned scheduler with a uniform $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ baseline using Llama 3.2-1B Instruct. The learned scheduler achieves slightly better average accuracy (%) (46.82 vs. 46.23), while additionally introducing stochastic $\lambda$ sampling as a regularization mechanism, improving robustness.